OpenStudy (anonymous):

A researcher found a significant relationship between a student’s IQ, a, the grade point average, b, and the score, y, on the verbal section of the SAT test. The relationship can be represented by the multiple regression equation y = 250 + 1.5a + 80b. Predict the SAT verbal score (to the nearest whole number) of a student whose IQ is 130 and grade point average is 2.2?

OpenStudy (ganpat):

given IQ = 130... i.e. a = 130 grade point = b = 2.2 Now substitute this in equation... y = 250 + 1.5a + 80b can u solve further ??

OpenStudy (anonymous):

this is already given in the question..

OpenStudy (anonymous):

I don't know how to...

OpenStudy (ganpat):

lol, that is the beauty of maths, everything is already given.. you just need to arrange in proper manner.. :D

OpenStudy (ganpat):

given IQ = 130... i.e. a = 130 grade point = b = 2.2 Now substitute this in equation... y = 250 + 1.5a + 80b y = 250 + 1.5 (130) + 80 (2.2) y = 621 is ma score... you recheck he calculations..

OpenStudy (anonymous):

hmmm let me check the possible answer given

OpenStudy (anonymous):

625, 621, 421and 371,.....your right.

OpenStudy (ganpat):

so, i have a good IQ :-p

OpenStudy (anonymous):

I didn't know that you have to multiple with the orignal equation .. I was replacing by the value given in the orignal question...stupid me lol

OpenStudy (ganpat):

now, next time onward, u be careful :).. ALL the Best !!

OpenStudy (anonymous):

Thank u I will...

OpenStudy (anonymous):

What do u think about the attached file..I already got 1 and 2